OpenStudy (anonymous):

Solve for the variable. y + -4 = -15 A. -19 B. -11 C. 19 D. 11

OpenStudy (saifoo.khan):

y = -15 - 4 y = -19

OpenStudy (precal):

y=-15+4 y=-11

OpenStudy (precal):

add 4 to both sides